OpenStudy (anonymous):

A 12-ft-by-15-ft rectangular swimming pool has a 3-ft-wide no-slip surface around it. What is the outer perimeter of the no-slip surface? (1 point) 78 ft 78 ft2 198 ft 198 ft2

OpenStudy (anonymous):

12+3+3 = 18,* 2 sides = 36 15+3+3= 21 * 2 sides = 42 add the 4 sides, 36+42 = 78

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

i got it :D hahaha